OpenStudy (anonymous):

Simplify the expression: sqrt (-6) and sqrt (-42)

OpenStudy (anonymous):

\[\sqrt{-6}+\sqrt{-42}=\sqrt{6}i+\sqrt{42}i\] there is really nothing you can do with this

OpenStudy (anonymous):

unless you want to write \[\sqrt{42}i=\sqrt{7}\sqrt{6}i\]

OpenStudy (anonymous):

I really don't know what this all means...Ive never been exposed to this in High School